OpenStudy (anonymous):

For the equation given below, evaluate y' at the point (−12). 5xy−4x+6=0

OpenStudy (anonymous):

implicit differentiation

OpenStudy (anonymous):

yeah i tried it and i get 5dy/dx =4 and i know i am suppose to divide by 5 but then i dont know where to plug in my point (-1,2)

OpenStudy (anonymous):

\[5xy'+5y-4=0\]

OpenStudy (anonymous):

you solve 5xy as a product

OpenStudy (anonymous):

oh so im suppose to do the product rule

OpenStudy (anonymous):

any time you differentiate x then y is like a constant

OpenStudy (anonymous):

when you differentiate y then x is like a constant followed by y'

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

my answer is different from yours

OpenStudy (anonymous):

absolutely\[Vdu+Udv\]

OpenStudy (anonymous):

NOW YOU PUT YOUR VALUES

OpenStudy (anonymous):

okay do you think you can check my answer i think i need to do it again

OpenStudy (anonymous):

\[5(-1)y'+5(2)-4=0\]

OpenStudy (anonymous):

BODMAS is important here

OpenStudy (anonymous):

\[-5y'+10-4=0\]

OpenStudy (anonymous):

so my final answer was 6/5 is that correct

OpenStudy (anonymous):

\[y'=\frac{ 6 }{ 5}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks